### v2

# Hybrid Type-II and Type-III seesaw model for the muon $g-2$ anomaly

Abstract
In this work we investigate the muon anomalous dipole moment $a_{\mu}$ in a model that extends the Standard Model with a scalar triplet and a lepton triplet. Different from previous studies, we find that there is still viable parameter space in this model to explain the discrepancy $\Delta a_{\mu}=a_{\mu}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})$. While being consistent with the current data of neutrino mass, electroweak precision measurements and the perturbativity of couplings, our model can provide new physics contribution $a_{\mu}^{\textrm{NP}}$ to cover the central region of $\Delta a_{\mu}$ with new scalar and lepton mass as low as around TeV. This mass scale is allowed by the current collider searches for doubly charged scalars and the lepton triplet, and they can be tested at future high energy and/or high luminosity colliders.
preprint: APCTP-Pre2024-017
I Introduction
The anomalous magnetic moment (AMM) of muon, denoted as $a_{\mu}\equiv(g-2)_{\mu}/2$, has been theoretically predicted in the Standard Model (SM) and experimentally measured both to very high precision. Since the $g-2$ experiment E821 performed at Brookhaven National Laboratory (BNL) released its data about two decades ago [1], a discrepancy $\Delta a_{\mu}=a_{\mu}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})$ has been existing and triggered rich phenomenological studies (see reviews [2, 3] and and references therein). Recently, on the theoretical side, a comprehensive summary of the most accurate SM prediction for $a_{\mu}$ is provided in [4] where the value is reported as $a_{\mu}(\mathrm{SM})=116591810(43)\times 10^{-11}$. On the experimental side, the Muon $g-2$ Experiment at Fermilab released its Run-1 dataset in 2021 and the result was $a_{\mu}^{2021}(\mathrm{Exp})=116592061(41)\times 10^{-11}$ [5] when combining the data from Fermilab Run-1 and BNL. This resulted in a deviation from the SM prediction of $a_{\mu}^{2021}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})=(251\pm 59)\times 10^{-11}$ with a significance of $4.2\,\sigma$. In 2023, Fermilab released data from Run-2 and Run-3 leading to a new result after being combined with Run-1 and the BNL data as $a_{\mu}^{2023}(\mathrm{Exp})=116592059(22)\times 10^{-11}$ [6]. In this case the deviation from the SM prediction $\Delta a_{\mu}=a_{\mu}^{2023}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})=(249\pm 48)\times 10^{-11}$ achieved a significance of $5.1\,\sigma$  [^1] Recently, there have been some discussions on the discrepancies in the calculation of the hadronic vacuum polarization contribution to $a_{\mu}(\mathrm{SM})$ between the lattice QCD calculations and experimental data in measurements of $e^{-}e^{+}\to\pi^{-}\pi^{+}$, which seems to reduce $\Delta a_{\mu}$ to some extend [7, 8, 9]. . These deviations inspired phenomenological investigations in various new physics models. Typical examples include the two-Higgs-doublet model [10, 11], the dark photon model [12], the supersymmetric models [13, 14, 15, 16, 17, 18, 14], leptoquark models [19, 20, 21], and vector-like lepton extended models [22, 23]. More references can be found in reviews [2, 3, 24, 25, 26, 27].
The chiral structure of the $a_{\mu}$ can be described by the following effective tensor operator
$\mathcal{L}^{\rm AMM}_{\rm eff.}=-\frac{e_{\ell}a_{\ell}}{4m_{\ell}}\,\overline{\ell_{L}}\,\sigma^{\mu\nu}\,\ell_{R}\,F_{\mu\nu}\,+{\rm h.c.}\,,$ (1)
in which the SM charged lepton $\ell$ should be understood as muon flavor. We can see that both the left-handed (LH) $\ell_{L}$ and the right-handed (RH) $\ell_{R}$ of muon are involved. If one considers the new physics contribution to $a_{\mu}$ at 1-loop level matching the required chiral structure, proper chiral flip is needed along the fermion line. This chiral flip can take place either in the external muon line realized by the SM Yukawa interaction outside the loop or in the internal fermion line inside the loop [26].
In the framework of simplified models, contributions to $\Delta a_{\mu}$ can be generated at the 1-loop level by introducing several new physics fields with various spins and quantum numbers under the electroweak gauge group ${\rm SU(2)_{L}\times U(1)_{Y}}$. As shown in [27], many models fail to explain $\Delta a_{\mu}$ either due to the wrong sign of predicted $\Delta a_{\mu}$ or excluded by other experimental constraints such as electroweak precision observables, flavor physics, and collider searches. Specifically, to our interest, it has been reported that simplified models introducing two new fields with different spins to the SM are unable to generate the chiral enhancement at the 1-loop level [27]. As a result, these models tend to confine the new physics particles to a low and compressed mass region in order to explain $\Delta a_{\mu}$ while avoiding collider constraints. However, an important consideration has been overlooked in the aforementioned analyses, i.e. the Yukawa interaction involving the SM Higgs doublet and the new physics fermion. The Yukawa interaction as a four-dimensional operator can naturally exist and induce mass mixing between the fermions in the SM and the new physics sector, thus providing an additional source of chiral flip to $a_{\mu}$.
In this work, we demonstrate that by incorporating the SM Higgs Yukawa interaction into the analysis, simplified models introducing two new physics fields to the SM can effectively explain $\Delta a_{\mu}$. As a concrete example, we consider a simplified model that extends the SM with a scalar triplet $(3,-1)$ and a lepton triplet $(3,0)$ under the electroweak gauge group ${\rm SU(2)_{L}\times U(1)_{Y}}$. These two fields have been scrutinized in neutrino mass generation mechanisms as Type-II [28, 29, 30, 31, 32, 33] and Type-III [34, 35] seesaw models, respectively. However, concentrating specifically on the physics of $(g-2)_{\mu}$, in this work we do not require our model to produce the experimentally suggested texture of neutrino mass matrix. Instead, we only ensure the theoretically predicted neutrino masses remains to be negligibly small.
This paper is organized as follows. In Sec. II we articulate our model setup and the related parameters. In Sec. III we present our analytical and numerical results. We draw our conclusion in Sec. IV.
II Model setup and parameters
II.1 Model setup
In our model we extend the SM with a scalar triplet $S$ and a left-handed lepton triplet $F_{L}$, of which the representation $(n,\,Y)$ under the SM electroweak gauge group ${\rm\mathrm{SU(2)_{L}}\times\mathrm{U(1)_{Y}}}$ are
[…]
and the component fields are
$S\equiv\Big{[}\begin{array}[]{cc}\delta^{-}/\sqrt{2}&\quad(v_{\delta}+\delta^{0}+i\,a^{0})/\sqrt{2}\\ \delta^{--}&-\delta^{-}/\sqrt{2}\end{array}\Big{]}\,,\qquad F_{L}\equiv\Big{[}\begin{array}[]{cc}F_{L}^{0}/\sqrt{2}&F_{L}^{+}\\ F_{L}^{-}&-F_{L}^{0}/\sqrt{2}\end{array}\Big{]}\,,$ (7)
in which $v_{\delta}$ is the vacuum expectation value (vev) of the neutral component of $S$ after electroweak symmetry breaking (EWSB). We consider the following mass and Yukawa terms in the Lagrangian which are most relevant to the physics of $(g-2)_{\mu}$
$\mathcal{L}_{\textrm{mass+Yuk.}}\supset$ $-\frac{1}{2}M_{F}\mathrm{Tr}\left[\overline{F_{L}}(F_{L})^{C}\right]-y^{ij}\overline{L_{L}^{i}}\ell_{R}^{j}H$ (8)
$-x_{L}^{ij}\overline{L_{L}^{i}}S\epsilon(L_{L}^{j})^{C}-\lambda_{L}^{i}\overline{\ell_{R}^{i}}\mathrm{Tr}\left[F_{L}S\right]-z_{L}^{i}\overline{L_{L}^{i}}(F_{L})^{C}\epsilon H^{\ast}+\mathrm{h.c.}\,,$
in which $H$ is the SM Higgs doublet and $\epsilon\equiv i\sigma^{2}$ is the antisymmetric tensor. $L^{i/j}_{L}\equiv(\nu^{i/j}_{L},\,\ell^{i/j}_{L})^{T}$ and $\ell^{i/j}_{R}$ are the LH doublet and RH singlet of SM lepton under ${\rm\mathrm{SU(2)_{L}}}$ with $i,j=1,2,3$ denoting the generation index, respectively. $(F_{L})^{C}$ denotes the charge conjugation of $F_{L}$ satisfying $(F_{L})^{C}\equiv C\overline{F_{L}}^{T}$. We require all Yukawa couplings in our model to be real in order to avoid constraints from CP-violating observables.
After EWSB we have the expansion $H=[G^{+},({v_{h}}+h+i\,G^{0})/\sqrt{2}]^{T}$ and can derive the following neutrino and charged lepton mass matrices
$\mathcal{L}_{\textrm{mass}}=$ $-\frac{1}{2}\left[\begin{array}[]{cc}\overline{\nu_{L}^{i}}&\overline{F_{L}^{0}}\end{array}\right]\left[\begin{array}[]{cc}-\sqrt{2}x_{L}^{ij}v_{\delta}&\frac{1}{2}z_{L}^{i}{v_{h}}\\ \frac{1}{2}z_{L}^{i}{v_{h}}&M_{F}\end{array}\right]\left[\begin{array}[]{cc}(\nu_{L}^{j})^{C}\\ (F_{L}^{0})^{C}\end{array}\right]$ (9)
$-\left[\begin{array}[]{cc}\overline{\ell_{L}^{i}}&\overline{F_{L}^{-}}\end{array}\right]\left[\begin{array}[]{cc}\frac{1}{\sqrt{2}}y^{ij}{v_{h}}&\frac{1}{\sqrt{2}}z_{L}^{i}{v_{h}}\\ \frac{1}{\sqrt{2}}(\lambda_{L}^{i})^{\ast}v_{\delta}&M_{F}\end{array}\right]\left[\begin{array}[]{cc}\ell_{R}^{j}\\ (F_{L}^{+})^{C}\end{array}\right]+\mathrm{h.c.}\,.$
Then the Yukawa interactions can be written as
$\mathcal{L}_{\textrm{Yuk.}}=$ $-\frac{1}{\sqrt{2}}y^{ij}h\overline{\ell_{L}^{i}}\ell_{R}^{j}-\frac{1}{\sqrt{2}}z_{L}^{i}h\overline{\ell_{L}^{i}}(F_{L}^{+})^{C}-\frac{1}{2}z_{L}^{i}h\overline{\nu_{L}^{i}}(F_{L}^{0})^{C}$ (10)
$-\frac{1}{\sqrt{2}}\lambda_{L}^{i}\delta^{0}\overline{\ell_{R}^{i}}F_{L}^{-}-\lambda_{L}^{i}\delta^{-}\overline{\ell_{R}^{i}}F_{L}^{0}-\lambda_{L}^{i}\delta^{--}\overline{\ell_{R}^{i}}F_{L}^{+}$
$+\frac{1}{\sqrt{2}}x_{L}^{ij}\delta^{0}\overline{\nu_{L}^{i}}(\nu_{L}^{j})^{C}-\frac{1}{\sqrt{2}}x_{L}^{ij}\delta^{-}\overline{\nu_{L}^{i}}(\ell_{L}^{j})^{C}$
$-\frac{1}{\sqrt{2}}x_{L}^{ij}\delta^{-}\overline{\ell_{L}^{i}}(\nu_{L}^{j})^{C}-x_{L}^{ij}\delta^{--}\overline{\ell_{L}^{i}}(\ell_{L}^{j})^{C}+\mathrm{h.c.}\,.$
In the above expression, the replacement $\delta^{0}\to i\,a^{0}$ would generate the Yukawa interactions of $a^{0}$ and we do not write them explicitly for compactness. Moreover, as will be discussed later, $G^{0}$ ($G^{\pm}$) will mix with $\delta^{0}$ ($\delta^{\pm}$) and yield mass eigenstate with zero mass, corresponding to the Goldstone bosons absorbed to be the longitudinal components of SM gauge bosons $Z$ ($W^{\pm}$) via the Higgs mechanism. Therefore, we also do not write terms of $G^{0},G^{\pm}$ explicitly for compactness.
For simplicity, we consider the scenario that only the second generation $i,j\equiv 2$ exist in Eq. (8), i.e. the mass mixing in Eqs. (9) and (10) would generate mass eigenstates of charged lepton and neutrino in SM for muon flavor, as well as heavy neutral and charged leptons in the new physics sector. Therefore, the indices $i,j$ will be dropped in the following. We can diagonalize the previous mass matrices through the following rotations
$\left[\begin{array}[]{c}\nu_{L}\\ F_{L}^{0}\end{array}\right]$ $\rightarrow\left[\begin{array}[]{cc}c_{L}^{\nu}&s_{L}^{\nu}\\ -s_{L}^{\nu}&c_{L}^{\nu}\end{array}\right]\left[\begin{array}[]{c}\nu_{L}\\ F_{L}^{0}\end{array}\right]\,,$ (11)
$\left[\begin{array}[]{c}\ell_{L}\\ F_{L}^{-}\end{array}\right]$ $\rightarrow\left[\begin{array}[]{cc}c_{L}^{\ell}&s_{L}^{\ell}\\ -s_{L}^{\ell}&c_{L}^{\ell}\end{array}\right]\left[\begin{array}[]{c}\ell_{L}\\ F_{L}^{-}\end{array}\right]\,,$
$\left[\begin{array}[]{c}\ell_{R}\\ (F_{L}^{+})^{C}\end{array}\right]$ $\rightarrow\left[\begin{array}[]{cc}c_{R}^{\ell}&s_{R}^{\ell}\\ -s_{R}^{\ell}&c_{R}^{\ell}\end{array}\right]\left[\begin{array}[]{c}\ell_{R}\\ (F_{L}^{+})^{C}\end{array}\right]\,.$
In the above, $s_{L/R}^{\nu/\ell}$ and $c_{L/R}^{\nu/\ell}$ are abbreviations of $\sin\theta_{L/R}^{\nu/\ell}$ and $\cos\theta_{L/R}^{\nu/\ell}$ when applicable. After the transformations in Eq. (11) we obtain the following Yukawa interactions in terms of mass eigenstates
$\mathcal{L}_{\textrm{Yuk.}}=\enspace\mathcal{L}_{\textrm{Yuk.}}^{\textrm{I}}+\mathcal{L}_{\textrm{Yuk.}}^{\textrm{II}}+\mathcal{L}_{\textrm{Yuk.}}^{\textrm{III}}\,,$ (12)
$\mathcal{L}_{\textrm{Yuk.}}^{\textrm{I}}=$ $\quad\,\overline{\ell_{L}}\ell_{R}[(-\frac{y}{\sqrt{2}}c_{L}^{\ell}c_{R}^{\ell}+\frac{z_{L}}{\sqrt{2}}c_{L}^{\ell}s_{R}^{\ell})h+\frac{1}{\sqrt{2}}\lambda_{L}^{\ast}s_{L}^{\ell}c_{R}^{\ell}\delta^{0}]$ (13)
$+\overline{\ell_{L}}(F_{L}^{+})^{C}[(-\frac{y}{\sqrt{2}}c_{L}^{\ell}s_{R}^{\ell}-\frac{z_{L}}{\sqrt{2}}c_{L}^{\ell}c_{R}^{\ell})h+\frac{1}{\sqrt{2}}\lambda_{L}^{\ast}s_{L}^{\ell}s_{R}^{\ell}\delta^{0}]$
$+\overline{F_{L}^{-}}\ell_{R}[(-\frac{y}{\sqrt{2}}s_{L}^{\ell}c_{R}^{\ell}+\frac{z_{L}}{\sqrt{2}}s_{L}^{\ell}s_{R}^{\ell})h-\frac{1}{\sqrt{2}}\lambda_{L}^{\ast}c_{L}^{\ell}c_{R}^{\ell}\delta^{0}]$
$+\overline{F_{L}^{-}}(F_{L}^{+})^{C}[(-\frac{y}{\sqrt{2}}s_{L}^{\ell}s_{R}^{\ell}-\frac{z_{L}}{\sqrt{2}}s_{L}^{\ell}c_{R}^{\ell})h-\frac{1}{\sqrt{2}}\lambda_{L}^{\ast}c_{L}^{\ell}s_{R}^{\ell}\delta^{0}]+\mathrm{h.c.}\,,$
$\mathcal{L}_{\textrm{Yuk.}}^{\textrm{II}}=$ $\quad\,\delta^{--}[-x_{L}(c_{L}^{\ell})^{2}\overline{\ell_{L}}(\ell_{L})^{C}+\lambda_{L}s_{R}^{\ell}c_{R}^{\ell}\overline{\ell_{R}}(\ell_{R})^{C}]$ (14)
$+\delta^{--}[-2x_{L}s_{L}^{\ell}c_{L}^{\ell}\overline{\ell_{L}}(F_{L}^{-})^{C}+\lambda_{L}\big{(}(s_{R}^{\ell})^{2}-(c_{R}^{\ell})^{2}\big{)}\overline{\ell_{R}}F_{L}^{+}]$
$+\delta^{--}[-x_{L}(s_{L}^{\ell})^{2}\overline{F_{L}^{-}}(F_{L}^{-})^{C}-\lambda_{L}s_{R}^{\ell}c_{R}^{\ell}\overline{(F_{L}^{+})^{C}}F_{L}^{+}]$
$+\delta^{-}[-\sqrt{2}x_{L}c_{L}^{\nu}c_{L}^{\ell}\overline{\nu_{L}}(\ell_{L})^{C}+\lambda_{L}s_{L}^{\nu}c_{R}^{\ell}\overline{\ell_{R}}\nu_{L}]$
$+\delta^{-}[-\sqrt{2}x_{L}c_{L}^{\nu}s_{L}^{\ell}\overline{\nu_{L}}(F_{L}^{-})^{C}-\lambda_{L}c_{L}^{\nu}c_{R}^{\ell}\overline{\ell_{R}}F_{L}^{0}]$
$+\delta^{-}[-\sqrt{2}x_{L}s_{L}^{\nu}c_{L}^{\ell}\overline{F_{L}^{0}}(\ell_{L})^{C}+\lambda_{L}s_{L}^{\nu}s_{R}^{\ell}\overline{(F_{L}^{+})^{C}}\nu_{L}]$
$+\delta^{-}[-\sqrt{2}x_{L}s_{L}^{\nu}s_{L}^{\ell}\overline{F_{L}^{0}}(F_{L}^{-})^{C}-\lambda_{L}c_{L}^{\nu}s_{R}^{\ell}\overline{(F_{L}^{+})^{C}}F_{L}^{0}]+\mathrm{h.c.}\,,$
$\mathcal{L}_{\textrm{Yuk.}}^{\textrm{III}}=$ $\quad\,\overline{\nu_{L}}(\nu_{L})^{C}[\frac{1}{2}z_{L}s_{L}^{\nu}c_{L}^{\nu}h+\frac{1}{\sqrt{2}}x_{L}(c_{L}^{\nu})^{2}\delta^{0}]$ (15)
$+\overline{\nu_{L}}(F_{L}^{0})^{C}[\frac{1}{2}z_{L}\big{(}(s_{L}^{\nu})^{2}-(c_{L}^{\nu})^{2}\big{)}h+\sqrt{2}x_{L}s_{L}^{\nu}c_{L}^{\nu}\delta^{0}]$
$+\overline{F_{L}^{0}}(F_{L}^{0})^{C}[-\frac{1}{2}z_{L}s_{L}^{\nu}c_{L}^{\nu}h+\frac{1}{\sqrt{2}}x_{L}(s_{L}^{\nu})^{2}\delta^{0}]+\mathrm{h.c.}\,.$
Then, the $(g-2)_{\mu}$ related Yukawa interactions are collected and simplified as
$\mathcal{L}_{\textrm{Yuk.}}=$ $\quad\,\overline{\ell_{L}}\ell_{R}[-\frac{m_{\ell}}{{v_{h}}}(c_{L}^{\ell})^{2}h+\frac{1}{\sqrt{2}}\lambda_{L}^{\ast}s_{L}^{\ell}c_{R}^{\ell}\delta^{0}]$ (16)
$+\overline{\ell_{L}}(F_{L}^{+})^{C}[-\frac{m_{F^{\pm}}}{{v_{h}}}s_{L}^{\ell}c_{L}^{\ell}h+\frac{1}{\sqrt{2}}\lambda_{L}^{\ast}s_{L}^{\ell}s_{R}^{\ell}\delta^{0}]$
$+\overline{F_{L}^{-}}\ell_{R}[-\frac{m_{\ell}}{{v_{h}}}s_{L}^{\ell}c_{L}^{\ell}h-\frac{1}{\sqrt{2}}\lambda_{L}^{\ast}c_{L}^{\ell}c_{R}^{\ell}\delta^{0}]$
$+\delta^{--}[-x_{L}(c_{L}^{\ell})^{2}\overline{\ell_{L}}(\ell_{L})^{C}+\lambda_{L}s_{R}^{\ell}c_{R}^{\ell}\overline{\ell_{R}}(\ell_{R})^{C}]$
$+\delta^{--}[-2x_{L}s_{L}^{\ell}c_{L}^{\ell}\overline{\ell_{L}}(F_{L}^{-})^{C}+\lambda_{L}\big{(}(s_{R}^{\ell})^{2}-(c_{R}^{\ell})^{2}\big{)}\overline{\ell_{R}}F_{L}^{+}]$
$+\delta^{-}[-\sqrt{2}x_{L}c_{L}^{\nu}c_{L}^{\ell}\overline{\nu_{L}}(\ell_{L})^{C}+\lambda_{L}s_{L}^{\nu}c_{R}^{\ell}\overline{\ell_{R}}\nu_{L}]$
$+\delta^{-}[-\lambda_{L}c_{L}^{\nu}c_{R}^{\ell}\overline{\ell_{R}}F_{L}^{0}-\sqrt{2}x_{L}s_{L}^{\nu}c_{L}^{\ell}\overline{F_{L}^{0}}(\ell_{L})^{C}]+\mathrm{h.c.}\,.$
Again, note that the replacement $\delta^{0}\to i\,a^{0}$ in $\mathcal{L}_{\textrm{Yuk.}}$ would generate the Yukawa interactions of $a^{0}$ and we do not write them explicitly for compactness.
As for the scalar sector, there can be rich interactions involving $H$ and $S$ [36] as follows
$V(H,S)=$ $-m_{H}^{2}H^{{\dagger}}H+\frac{\lambda}{4}(H^{{\dagger}}H)^{2}+M_{S}^{2}\mathrm{Tr}[S^{{\dagger}}S]+[\mu_{HS}(H^{T}\epsilon SH)+\mathrm{h.c.}]$
$+\lambda_{1}(H^{{\dagger}}H)\mathrm{Tr}[S^{{\dagger}}S]+\lambda_{2}(\mathrm{Tr}[S^{{\dagger}}S])^{2}+\lambda_{3}\mathrm{Tr}[(S^{{\dagger}}S)^{2}]+\lambda_{4}H^{{\dagger}}S^{{\dagger}}SH\,,$ (17)
which would generate the mixing between $H$ and $S$ and result in the mass eigenstates including the SM Higgs boson and several scalars in the new physics sector. Similar to Eq. (11) we can diagnolize the electrically neutral CP-even scalar mass matrix by performing the following rotation
$\left[\begin{array}[]{c}h\\ \delta^{0}\end{array}\right]$ $\rightarrow\left[\begin{array}[]{cc}c^{h}&s^{h}\\ -s^{h}&c^{h}\end{array}\right]\left[\begin{array}[]{c}h\\ \delta^{0}\end{array}\right]\,,$ (18)
$\left[\begin{array}[]{c}G^{0}\\ a^{0}\end{array}\right]$ $\rightarrow\left[\begin{array}[]{cc}c^{a}&s^{a}\\ -s^{a}&c^{a}\end{array}\right]\left[\begin{array}[]{c}G^{0}\\ a^{0}\end{array}\right]\,,$
$\left[\begin{array}[]{c}(G^{+})^{\ast}\\ \delta^{-}\end{array}\right]$ $\rightarrow\left[\begin{array}[]{cc}c^{G}&s^{G}\\ -s^{G}&c^{G}\end{array}\right]\left[\begin{array}[]{c}(G^{+})^{\ast}\\ \delta^{-}\end{array}\right]\,,$
in which $s^{h/a/G},c^{h/a/G}$ are the abbreviations of $\sin\theta^{h/a/G},\cos\theta^{h/a/G}$ following the same convention as Eq. (11), and $G^{0},G^{\pm}$ are the Goldstone bosons which will be absorbed to be the longitudinal components of SM gauge bosons $Z,W^{\pm}$ via the Higgs mechanism. For later analysis we will label the masses of physical scalars in the mass eigenstates as $m_{h},m_{\delta^{0}},m_{a^{0}},m_{\delta^{-}},m_{\delta^{--}}$, with $h,\delta^{0}$ being the neutral CP-even scalars, $a^{0}$ being the neutral CP-odd scalar, and $\delta^{-},\delta^{--}$ being the singly and doubly charged scalars, respectively. Details of the diagnolization and rotation matrices in the scalar sector can be found in e.g. [36].
II.2 Relations of parameters
When diagnolizing the neutrino mass matrix in Eq. (9), we have the following relations
$-\sqrt{2}x_{L}v_{\delta}$ $=m_{\nu}(c_{L}^{\nu})^{2}+m_{F^{0}}(s_{L}^{\nu})^{2}\,,$
$z_{L}{v_{h}}$ $=2(m_{F^{0}}-m_{\nu})s_{L}^{\nu}c_{L}^{\nu}\,,$
$M_{F}$ $=m_{\nu}(s_{L}^{\nu})^{2}+m_{F^{0}}(c_{L}^{\nu})^{2}\,.$ (19)
Similarly, when diagnolizing the charged lepton mass matrix in Eq. (9), we have the following relations
$\frac{1}{\sqrt{2}}y{v_{h}}$ $=m_{\ell}c_{L}^{\ell}c_{R}^{\ell}+m_{F^{\pm}}s_{L}^{\ell}s_{R}^{\ell}\,,\quad\quad\frac{1}{\sqrt{2}}z_{L}{v_{h}}=-m_{\ell}c_{L}^{\ell}s_{R}^{\ell}+m_{F^{\pm}}s_{L}^{\ell}c_{R}^{\ell}\,,$
$\frac{1}{\sqrt{2}}\lambda_{L}^{\ast}v_{\delta}$ $=-m_{\ell}s_{L}^{\ell}c_{R}^{\ell}+m_{F^{\pm}}c_{L}^{\ell}s_{R}^{\ell}\,,\quad\quad\quad\,\,M_{F}=m_{\ell}s_{L}^{\ell}s_{R}^{\ell}+m_{F^{\pm}}c_{L}^{\ell}c_{R}^{\ell}\,.$ (20)
Given that there are different equalities on $z_{L}v_{h}$ and $M_{F}$ in Eqs. (II.2) and (II.2), we can have the following identities
$m_{F^{0}}$ $=\frac{1}{(c_{L}^{\nu})^{2}}\left(m_{\ell}s_{L}^{\ell}s_{R}^{\ell}+m_{F^{\pm}}c_{L}^{\ell}c_{R}^{\ell}-m_{\nu}(s_{L}^{\nu})^{2}\right),$ (21)
$-m_{\ell}c_{L}^{\ell}s_{R}^{\ell}+m_{F^{\pm}}s_{L}^{\ell}c_{R}^{\ell}$ $=\sqrt{2}\frac{s_{L}^{\nu}}{c_{L}^{\nu}}\left(m_{\ell}s_{L}^{\ell}s_{R}^{\ell}+m_{F^{\pm}}c_{L}^{\ell}c_{R}^{\ell}-m_{\nu}\right)\,.$
Considering that the current status of SM neutrino mass measurements from cosmology suggest $\sum_{\nu}m_{\nu}<1\,\textrm{eV}$ [37, 38], we would apply $m_{\nu}\to 0$ as a constraint on the model parameters in the above relations. As for the SM charged lepton masses $m_{\ell}$ we utilize the non-zero value provided by Particle Data Group [38] in numerical calculations. However, one can still apply $m_{\ell}\to 0$ to obtain more compact analytical relations. For Eq. (II.2) we have
$-\sqrt{2}x_{L}v_{\delta}\approx m_{F^{0}}(s_{L}^{\nu})^{2},\quad z_{L}{v_{h}}\approx 2m_{F^{0}}s_{L}^{\nu}c_{L}^{\nu},\quad M_{F}\approx m_{F^{0}}(c_{L}^{\nu})^{2}\,,$ (22)
and for Eq. (II.2) we have
$\frac{1}{\sqrt{2}}y{v_{h}}$ $\approx m_{F^{\pm}}s_{L}^{\ell}s_{R}^{\ell}\,,\quad\quad\,\,\frac{1}{\sqrt{2}}z_{L}{v_{h}}\approx m_{F^{\pm}}s_{L}^{\ell}c_{R}^{\ell}\,,$
$\frac{1}{\sqrt{2}}\lambda_{L}^{\ast}v_{\delta}$ $\approx m_{F^{\pm}}c_{L}^{\ell}s_{R}^{\ell}\,,\quad\quad\quad\quad\,\,M_{F}\approx m_{F^{\pm}}c_{L}^{\ell}c_{R}^{\ell}\,.$ (23)
Similarly, applying $m_{\nu}\rightarrow 0$ and $m_{\ell}\to 0$ would reduce Eq. (21) to the following simple form
$m_{F^{0}}\approx m_{F^{\pm}}\frac{c_{L}^{\ell}c_{R}^{\ell}}{(c_{L}^{\nu})^{2}}\,,\quad\tan\theta_{L}^{\ell}\approx\sqrt{2}\tan\theta_{L}^{\nu}\,.$ (24)
Based on the above discussion we have the following consideration on parameter setup.
  • $s_{L}^{\nu}$ characterizing the mixing of SM neutrino $\nu$ with $F^{0}$ is constrained by the electroweak precision measurements, for the muon flavor to be [39],
    $s^{\nu}_{L}\leq 0.017\,,$ (25)
    which can yield the following approximations with small mixing angle in the LH lepton sector
    $s_{L}^{\ell}\approx\sqrt{2}\,s_{L}^{\nu}\sim\mathcal{O}(10^{-2})\,,\quad c^{\nu}_{L}\approx c^{\ell}_{L}\approx 1\,,\quad m_{F^{0}}\approx m_{F^{\pm}}\,c^{\ell}_{R}\,.$ (26)
    This implies that nearly linear correlation exist between the above physical quantities which should be took into consideration when choosing independent input parameters.
  • $s_{R}^{\ell}$ characterizing the mixing of RH component of SM charged lepton $\ell$ with $F^{-}$ can be solved from Eqs. (22) and (II.2). In the parameter regions chosen for our numerical calculation it turns out to be the following  [^2] Note that in other models the suppression of $s^{\ell}_{R}$ compared to $s^{\ell}_{L}$ can be different from our model. Taking Type-III seesaw model as an example, the suppression is $\sim m_{\ell}/M_{F}$ [39, 40].
    $s^{\ell}_{R}\sim\mathcal{O}(10^{-1})\,s^{\ell}_{L}\,,$ (27)
    which means that the small angle approximation also holds for $\theta_{R}$ (see Eqs. (II.3) and (II.3) for more details). Therefore, in our model we have
    $s^{\ell}_{R}\sim\mathcal{O}(10^{-3})\,,\quad c^{\ell}_{R}\approx 1\,,\quad m_{F^{0}}\approx m_{F^{\pm}}\,.$ (28)
  • As the mass gap $\Delta m_{F}\equiv m_{F^{\pm}}-m_{F^{0}}$ in Eq. (24) is negligibly small, $\Delta m_{F}$ does not alter the main decay signals of the heavy leptons $F^{\pm},F^{0}$ as discussed in [41, 42]. Therefore, we can simply impose the similar lower bound of mass to be
    $m_{F^{\pm}}\,,\,m_{F^{0}}\,\gtrsim 1000\,{\rm GeV}\,.$ (29)
II.3 Input parameters
Now we determine the physically reasonable choice of input parameters of our model.
In the scalar sector shown in Eq. (II.1), despite the rich parameters and phenomenology about $V(H,S)$ (see e.g. [43] for collider signal searches), the physical quantities most relevant to $(g-2)_{\mu}$ are
$v_{\delta}\,,\,v_{h}\,,\qquad m_{h}\,,\,m_{\delta^{0}}\,,\,m_{a^{0}}\,,\,m_{\delta^{-}}\,,\,m_{\delta^{--}}\,,\qquad s^{h}\,,\,s^{a}\,,\,s^{G}\,,$ (30)
which are internally related [36]. Our requirements and simplifications include the following ones.
  • As for the two vevs $v_{h},v_{\delta}$, we require them to satisfy [44, 45]
    $v=\sqrt{v_{h}^{2}+2v_{\delta}^{2}}\approx 246\,{\rm GeV}\,,$ (31)
    to be consistent with the electroweak precision measurement [38] and
    $m_{h}=125\,{\rm GeV}\,,$ (32)
    to match the mass of the SM-like Higgs boson discovered at the Large Hadron Collider (LHC) [46, 47, 38]. To avoid the constraints $v_{\delta}\lesssim 5\,{\rm GeV}$ from the electroweak precision observables [48], in our numerical calculation we choose
    $v_{\delta}=5\,{\rm GeV}\,,\,v_{h}\approx 246\,{\rm GeV}\,.$ (33)
  • As for the masses of heavy scalars in the new physics sector, we require that the following approximations hold well
    $M_{S}\approx m_{\delta^{0}}\approx m_{a^{0}}\approx m_{\delta^{-}}\approx m_{\delta^{--}}\,,$ (34)
    which can be properly realized under the conditions of $v_{\delta}\ll v_{h}$ and $m_{h}\ll M_{S}$ [36]. Hereafter we would use $m_{\delta}$ as the notation to denote
    $m_{\delta}\equiv m_{\delta^{0}}\approx m_{a^{0}}\approx m_{\delta^{-}}\approx m_{\delta^{--}}\,.$ (35)
    Considering that the current searches for doubly charged scalar sets a lower limit of mass to be around $1\,{\rm TeV}$ assuming decaying to SM leptons [49, 50], we take the following benchmark throughout this work
    $m_{\delta}=1000\,{\rm GeV}\,.$ (36)
  • As for the mixing $s^{h},s^{a},s^{G}$ in the scalar sector, we first have the following requirement to safely pass the current constraints from the study of Higgs data [51, 52]
    $|s^{h}|\lesssim\mathcal{O}(0.1)\,,$ (37)
    which can be properly realized due to the rich parameter space of $V(H,S)$. Moreover, our choice of $v_{\delta},v_{h}$ in Eq. (31) and (33) also results in [36]
    $s^{a}\approx\tan\theta^{a}=\frac{2\,v_{\delta}}{v_{h}}\sim\mathcal{O}(10^{-2})\,,\quad s^{G}\approx\tan\theta^{G}=\frac{1}{\sqrt{2}}\tan\theta^{a}\sim\mathcal{O}(10^{-2})\,,$ (38)
    which make the mixing $s^{h},s^{a},s^{G}$ all in the small value region. While keeping $s^{h},s^{a},s^{G}$ in the analytical results, we will simply impose the following simplifications in numerical calculation which make negligible difference to the results
    $s^{h}=s^{a}=s^{G}=0\,.$ (39)
In the fermion sector shown in Eq. (8), we can convert the five parameters in the Lagrangian
$M_{F}\,,\quad y\,,\quad x_{L}\,,\quad\lambda_{L}\,,\quad z_{L}\,,$ (40)
to physical quantities in terms of mass eigenstates as follows
[…]
Despite there are seven quantities listed above, two of them are not independent as shown in Eq. (24). As for the five independent parameters, three of them can be naturally chosen as
$m_{\nu}\,,\quad\,m_{\ell}\,,\,\quad m_{F^{0}}\,.$ (42)
According to Eq. (28) and similar to Eq. (35), hereafter we would use $m_{F}$ as the following notation
$m_{F}\equiv m_{F^{0}}\approx m_{F^{\pm}}\,,$ (43)
to denote the numerical approximations, while keeping in mind they can be independent quantities in principle as discussed above. As discussed in Eq. (29), in this work we would take
$m_{F}\gtrsim 1000\,{\rm GeV}\,.$ (44)
As for the choice of the other two independent parameters, given the almost linear correlation between $s^{\nu}_{L}$ and $s^{\ell}_{L}$ shown in Eq. (26), we should not choose them simultaneously. In this work, we consider the following two different schemes.
  • $\{\lambda_{L},x_{L}\}$ scheme
    In this scheme, Eqs. (22) and (II.2) suggest that the other parameters can be expressed as
    $z_{L}=2\sqrt{-\sqrt{2}\,x_{L}\frac{v_{\delta}\,m_{F}}{v_{h}^{2}}}\,,\quad y\approx 2\,\lambda_{L}\frac{v_{\delta}}{v_{h}}\sqrt{\frac{-x_{L}\,v_{\delta}}{\sqrt{2}\,m_{F}}}\,,$
    $s^{\nu}_{L}\approx\sqrt{-\sqrt{2}\,x_{L}\frac{v_{\delta}}{m_{F}}}\,,\qquad\,s^{\ell}_{L}\approx\sqrt{2}\,s^{\nu}_{L}\,,\qquad s^{\ell}_{R}\approx\frac{\lambda_{L}\,v_{\delta}}{\sqrt{2}\,m_{F}}\,.$ (45)
    Again, note that in our numerical analysis we choose all Yukawa couplings to be real and in our conventions we have $x_{L}<0$ and $y,z_{L},\lambda_{L}>0$.
  • $\{\lambda_{L},\theta_{L}^{\nu}\}$ scheme
    In this scheme, Eqs. (22) and (II.2) suggest that the other parameters can be expressed as
    $x_{L}\approx-\frac{1}{\sqrt{2}}(s^{\nu}_{L})^{2}\frac{m_{F}}{v_{\delta}}\,,\quad z_{L}\approx 2\,s^{\nu}_{L}\frac{m_{F}}{v_{h}}\,,\quad y\approx\sqrt{2}\,s^{\nu}_{L}\,\lambda_{L}\frac{v_{\delta}}{v_{h}}\,,$
    $s^{\ell}_{L}\approx\sqrt{2}\,s^{\nu}_{L}\,,\quad\quad\quad\quad\quad s^{\ell}_{R}\approx\frac{\lambda_{L}\,v_{\delta}}{\sqrt{2}\,m_{F}}\,.$ (46)
As we will discuss later, the $\{\lambda_{L},x_{L}\}$ and $\{\lambda_{L},\theta_{L}^{\nu}\}$ schemes are physically suitable for the illustration of the decoupling behavior and chiral enhancement effects in the predictions of $(g-2)_{\mu}$ in our model, respectively.
To sum up the above discussion, the input parameters in our analysis are arranged as
[…]
$\quad v_{\delta}=5\,\textrm{GeV}\,,\quad m_{\delta}=1000\,\textrm{GeV}\,,\quad\quad s^{h}=s^{a}=s^{G}=0\,,$
[…]
II.4 Perturbativity requirement
Based on the discussion in Section II.3 it is easy to study the perturbativity behavior of the Yukawa couplings in our model.
  • In the $\{\lambda_{L},x_{L}\}$ scheme, we have the following approximations from Eq. (II.3)
    $z_{L}\sim\sqrt{\frac{-x_{L}}{10^{4}}\,\frac{m_{F}}{\rm GeV}}\,,\quad y\sim\sqrt{\frac{-x_{L}}{100}\,\frac{\rm GeV}{m_{F}}}\,.$ (48)
    In this work we focus on $1\,{\rm TeV}\lesssim m_{F}\lesssim 5\,{\rm TeV}$ and $|x_{L}|\lesssim\mathcal{O}(10^{-1})$, thus the requirement of perturbativity $z_{L},|x_{L}|,y<\mathcal{O}(1)$ can be easily satisfied. Note that Eq. (25) and Eq. (II.3) also imply a lower bound of $m_{F}$ satisfying
    $m_{F}\gtrsim(-x_{L})\frac{\sqrt{2}\,v_{\delta}}{(0.017)^{2}}\approx 25\,|x_{L}|\,{\rm TeV}\,,$ (49)
    which will be manifested in our numerical results discussed later (see e.g. Fig. 2).
  • In the $\{\lambda_{L},\theta_{L}^{\nu}\}$ scheme, we have the following approximations from Eq. (II.3)
    $z_{L}\sim\frac{1}{100}\,\frac{m_{F}}{\rm GeV}\,s^{\nu}_{L}\,,\quad x_{L}\sim-\frac{1}{10}\,\frac{m_{F}}{\rm GeV}(s^{\nu}_{L})^{2}\,,\quad y\sim\frac{1}{10}s^{\nu}_{L}\,.$ (50)
    Given $s^{\nu}_{L}\leq 0.017$ indicated in Eq. (25), we can see that the requirement of perturbativity $z_{L},|x_{L}|,y<\mathcal{O}(1)$ can also be easily satisfied for the mass region $1\,{\rm TeV}\lesssim m_{F}\lesssim 5\,{\rm TeV}$ in our discussion.
III Analytical and numerical results
In this section we present our main results. We cross check our formulae by implementing our model in Eq. (8) to FeynRules [53, 54] interfaced to FeynArts [55] and FormCalc [56] to perform loop calculations. Then we extract from the amplitude to obtain the expressions of $a^{\textrm{NP}}_{\mu}$, i.e. the new physics contribution to $\Delta a_{\mu}$ in our model, and further reduce the loop functions to simple expressions via Package-X [57, 58].
III.1 Analytical results
Figure 1: Feynman diagrams contributing to the $(g-2)_{\mu}$ in which the left, middle and right panel corresponds to $a_{\mu}^{\ell,\,\textrm{total}},\,a_{\mu}^{F,\,\textrm{total}},\,a_{\mu}^{\nu,\,\textrm{total}}$ in Eq. (51), respectively. Note that in the left and middle diagrams the photon can be emitted from either of the two charged particles in the loop, while the emission only takes place from $\delta^{-}$ in the right diagram. The Feynman diagrams are drawn by JaxoDraw [59].
In Fig. 1 we show the Feynman diagrams contributing to $(g-2)_{\mu}$ in our model, in which the leptons $\ell,\nu$ should be understood to carry the muon flavor as $\mu,\nu_{\mu}$. They are divided into three parts according to the fermions appearing in the loop. Originating from the left, middle and right panel, respectively, the analytical results of $a^{\textrm{NP}}$ generated by our model can be decomposed to
$a^{\textrm{NP}}_{\mu}\equiv a_{\mu}^{\ell,\,\textrm{total}}+a_{\mu}^{F,\,\textrm{total}}+a_{\mu}^{\nu,\,\textrm{total}}\,.$ (51)
Note that all Yukawa couplings in our model are chosen to be real, but we will present our analytical results in the context of complex Yukawa couplings for the more general scenario. To write down the analytical results, we first define the following expressions as the reduced form of loop functions (see [60] for a more complete list).
  • In calculating $a_{\mu}^{\ell,\,\textrm{total}}$ from the left panel of Fig. 1 which satisfies $m_{\ell}\ll m_{h},\,m_{\delta}$, we define
    $F_{LL}^{f,1}(x)$ $=\frac{1}{6}+x(\frac{1}{2}\log x+\frac{25}{24})\,,\quad F_{LR}^{f,1}(x)=-\frac{1}{2}\log x-\frac{3}{4}+x(-2\log x-\frac{8}{3})\,,$
    $F_{LL}^{S,1}(x)$ $=-\frac{1}{12}+\frac{1}{8}x\,,\qquad\qquad\quad\,F_{LR}^{S,1}(x)=-\frac{1}{4}+x(-\frac{1}{2}\log x-\frac{11}{12})\,.$ (52)
  • In calculating $a_{\mu}^{F,\,\textrm{total}}$ from the middle panel of Fig. 1 which satisfies $m_{\ell}\ll m_{F},\,m_{\delta}$, we define
    $F_{LL}^{f,2}(x)$ $=\frac{2+3x-6x^{2}+x^{3}+6x\log x}{12(1-x)^{4}}\,,\quad\quad\,\,\,\,F_{LR}^{f,2}(x)=\frac{-3+4x-x^{2}-2\log x}{4(1-x)^{3}}\,,$
    $F_{LL}^{S,2}(x)$ $=-\frac{1-6x+3x^{2}+2x^{3}-6x^{2}\log x}{12(1-x)^{4}}\,,\quad F_{LR}^{S,2}(x)=\frac{-1+x^{2}-2x\log x}{4(1-x)^{3}}\,.$ (53)
  • In calculating $a_{\mu}^{\nu,\,\textrm{total}}$ from the right panel of Fig. 1 which satisfies $m_{\nu}\ll m_{\ell}\ll m_{\delta}$, we define
    $F_{LL}^{S,3}(x)=-\frac{1}{12}(1+\frac{1}{2}x)\,,\quad F_{LR}^{S,3}(x)=-\frac{1}{4}(1+\frac{2}{3}x)\,.$ (54)
The explicit form of $a_{\mu}^{\ell,\,\textrm{total}}$ originating from the left panel of Fig. 1 is
$a_{\mu}^{\ell,\,\textrm{total}}=a_{\mu}^{\ell,\,h}+a_{\mu}^{\ell,\,\delta^{0}}+a_{\mu}^{\ell,\,a^{0}}+a_{\mu}^{\ell,\,\delta^{--}}\,,$ (55)
in which the neutral CP-even scalars $h,\delta^{0}$ contribute as
$a_{\mu}^{\ell,\,h}$ $=\quad\frac{m_{\mu}^{2}}{4\pi^{2}m_{h}^{2}}[|\frac{m_{\mu}}{v_{h}}(c_{L}^{\ell})^{2}c^{h}+\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}c_{R}^{\ell}s^{h}|^{2}-\frac{m_{\mu}^{2}}{v^{2}}]\cdot F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{h}^{2}})$
$\quad+\frac{m_{\mu}^{2}}{4\pi^{2}m_{h}^{2}}\textrm{Re}[(\frac{m_{\mu}}{v_{h}}(c_{L}^{\ell})^{2}c^{h}+\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}c_{R}^{\ell}s^{h})^{2}-\frac{m_{\mu}^{2}}{v^{2}}]\cdot F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{h}^{2}})\,,$
$a_{\mu}^{\ell,\,\delta^{0}}$ $=\quad\frac{m_{\mu}^{2}}{4\pi^{2}m_{\delta^{0}}^{2}}|\frac{m_{\mu}}{v_{h}}(c_{L}^{\ell})^{2}s^{h}-\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}c_{R}^{\ell}c^{h}|^{2}\cdot F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{0}}^{2}})$
$\quad+\frac{m_{\mu}^{2}}{4\pi^{2}m_{\delta^{0}}^{2}}\textrm{Re}[(\frac{m_{\mu}}{v_{h}}(c_{L}^{\ell})^{2}s^{h}-\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}c_{R}^{\ell}c^{h})^{2}]\cdot F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{0}}^{2}})\,,$ (56)
and the neutral CP-odd scalar $a^{0}$ contributes as
$a_{\mu}^{\ell,\,a^{0}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{a^{0}}^{2}}|\lambda_{L}|^{2}(c^{a}s^{\ell}_{L}c^{\ell}_{R})^{2}\cdot F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{a^{0}}^{2}})$
$\quad+\frac{m_{\mu}^{2}}{8\pi^{2}m_{a^{0}}^{2}}\textrm{Re}[-(\lambda_{L}c^{a}s^{\ell}_{L}c^{\ell}_{R})^{2}]\cdot F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{a^{0}}^{2}})\,,$ (57)
and the charged scalar $\delta^{--}$ contributes as
$a_{\mu}^{\ell,\,\delta^{--}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{--}}^{2}}[|x_{L}|^{2}(2c_{L}^{\ell})^{2}+|\lambda_{L}|^{2}(2s_{R}^{\ell}c_{R}^{\ell})^{2}]\cdot[-F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})+2F_{LL}^{S,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})]$
$\quad+\frac{m_{\mu}^{2}}{4\pi^{2}m_{\delta^{--}}^{2}}\textrm{Re}[\big{(}-2x_{L}(c_{L}^{\ell})^{2}\big{)}(2\lambda_{L}^{\ast}s_{R}^{\ell}c_{R}^{\ell})]\cdot[-F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})]\,.$ (58)
Note that we have subtracted the SM Higgs contribution from $a_{\mu}^{\ell,\,h}$ to meet the definition of $a^{\textrm{NP}}$, i.e. the contribution exclusively generated from new physics sector. Moreover, the symmetry factor from the coupling of ¯$\bar{\ell}\ell^{C}\delta^{--}$ has been properly considered in $a_{\mu}^{\ell,\,\delta^{--}}$ as pointed out in Ref. [61].
The explicit form of $a_{\mu}^{F}$ originating from the middle panel of Fig. 1 is
$a_{\mu}^{F,\,\textrm{total}}=a_{\mu}^{F,\,h}+a_{\mu}^{F,\,\delta^{0}}+a_{\mu}^{F,\,a^{0}}+a_{\mu}^{F,\,\delta^{--}}+a_{\mu}^{F,\,\delta^{-}}\,,$ (59)
in which the neutral CP-even scalars $h,\delta^{0}$ contribute as
$a_{\mu}^{F,\,h}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{h}^{2}}[|\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}+\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}s^{h}|^{2}+|\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}-\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}s^{h}|^{2}]\cdot F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{h}^{2}})$
$\quad+\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{h}^{2}}\textrm{Re}[(\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}+\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}s^{h})(\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}-\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}s^{h})]\cdot F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{h}^{2}})\,,$
$a_{\mu}^{F,\,\delta^{0}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{0}}^{2}}[|\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}-\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}c^{h}|^{2}+|\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}+\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}c^{h}|^{2}]\cdot F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{0}}^{2}})$
$\quad+\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{0}}^{2}}\textrm{Re}[(\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}-\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}c^{h})(\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}+\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}c^{h})]\cdot F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{0}}^{2}})\,,$ (60)
and the neutral CP-odd scalar $a^{0}$ contributes as
$a_{\mu}^{F,\,a^{0}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{a^{0}}^{2}}[\frac{1}{2}|\lambda_{L}|^{2}(c^{a}s^{\ell}_{L}s^{\ell}_{R})^{2}+\frac{1}{2}|\lambda_{L}|^{2}(c^{a}c^{\ell}_{L}c^{\ell}_{R})^{2}]\cdot F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{a^{0}}^{2}})$
$\quad+\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{a^{0}}^{2}}\textrm{Re}[(\frac{1}{\sqrt{2}}\lambda_{L}c^{a}s^{\ell}_{L}s^{\ell}_{R})(\frac{1}{\sqrt{2}}\lambda_{L}c^{a}c^{\ell}_{L}c^{\ell}_{R})]\cdot F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{a^{0}}^{2}})\,,$ (61)
and the charged scalars $\delta^{-},\delta^{--}$ contribute as
$a_{\mu}^{F,\,\delta^{--}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{--}}^{2}}[4|x_{L}|^{2}(s_{L}^{\ell}c_{L}^{\ell})^{2}+|\lambda_{L}|^{2}\big{(}(c_{R}^{\ell})^{2}-(s_{R}^{\ell})^{2}\big{)}^{2}]\cdot[-F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LL}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]$
$\quad+\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{--}}^{2}}\textrm{Re}[(2x_{L}s_{L}^{\ell}c_{L}^{\ell})\big{(}\lambda_{L}^{\ast}((c_{R}^{\ell})^{2}-(s_{R}^{\ell})^{2})\big{)}]\cdot[-F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]\,,$
$a_{\mu}^{F,\,\delta^{-}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{-}}^{2}}[2|x_{L}|^{2}(s_{L}^{\nu}c_{L}^{\ell}c^{G})^{2}+|\lambda_{L}|^{2}(c_{L}^{\nu}c_{R}^{\ell}c^{G})^{2}]\cdot F_{LL}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})$
$\quad+\frac{m_{\mu}m_{F^{0}}}{4\pi^{2}m_{\delta^{-}}^{2}}\textrm{Re}[(\sqrt{2}x_{L}s_{L}^{\nu}c_{L}^{\ell}c^{G})(\lambda_{L}^{\ast}c_{L}^{\nu}c_{R}^{\ell}c^{G})]\cdot F_{LR}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})\,,$ (62)
The explicit form of $a_{\mu}^{\nu}$ originating from the right panel of Fig. 1 is
$a_{\mu}^{\nu,\,\textrm{total}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{-}}^{2}}[2|x_{L}|^{2}(c_{L}^{\nu}c_{L}^{\ell}c^{G})^{2}+|\lambda_{L}|^{2}(s_{L}^{\nu}c_{R}^{\ell}c^{G})^{2}]\cdot F_{LL}^{S,3}(\frac{m_{\mu}^{2}}{m_{\delta^{-}}^{2}})$
$\quad+\frac{m_{\mu}m_{\nu}}{4\pi^{2}m_{\delta^{-}}^{2}}\textrm{Re}[(-\sqrt{2}x_{L}c_{L}^{\nu}c_{L}^{\ell}c^{G})(\lambda_{L}^{\ast}s_{L}^{\nu}c_{R}^{\ell}c^{G})]\cdot F_{LR}^{S,3}(\frac{m_{\mu}^{2}}{m_{\delta^{-}}^{2}})\,.$ (63)
One can clearly see that for each one consisting $a_{\mu}^{F,\,\textrm{total}}$ in Eq. (59), the chiral enhancement appears as indicated by $m_{F^{0}}$ or $m_{F^{\pm}}$ in the numerator of the second line. As a result of this enhancement, $a_{\mu}^{F,\,\textrm{total}}$ in Eq. (51) turns out to play the absolutely dominant role compared to $a_{\mu}^{\ell,\,\textrm{total}}$ and $a_{\mu}^{\nu,\,\textrm{total}}$, i.e.
$a_{\mu}^{F,\,\textrm{total}}\,\gg\,a_{\mu}^{\ell,\,\textrm{total}}\,,\,a_{\mu}^{\nu,\,\textrm{total}}\,.$ (64)
Furthermore, for the different contributions to $a_{\mu}^{F,\,\textrm{total}}$ in Eq. (59) , we find that $a_{\mu}^{F,\,\delta^{--}}$ and $a_{\mu}^{F,\,\delta^{-}}$ contribute the dominant and sub-dominant part, respectively.
To simplify our analytical results and highlight the chiral enhancement, we utilize the small mixing condition $s^{\nu}_{L},s^{\ell}_{L},s^{\ell}_{R},s^{h},s^{a},s^{G}\ll 1$ discussed in Section II.2 and II.3 and extract the most relevant terms in Eq. (59) as follows
$a^{\textrm{NP}}_{\mu}\approx$ $\,a_{\mu}^{F,\,\textrm{total}}\approx\,a_{\mu}^{F,\,\delta^{--}}+a_{\mu}^{F,\,\delta^{-}}\,,$
$\approx$ $\quad\quad\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{--}}^{2}}\textrm{Re}[(2x_{L}s_{L}^{\ell}c_{L}^{\ell})\big{(}\lambda_{L}^{\ast}((c_{R}^{\ell})^{2}-(s_{R}^{\ell})^{2})\big{)}]\cdot[-F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]$
$+\frac{m_{\mu}m_{F^{0}}}{4\pi^{2}m_{\delta^{-}}^{2}}\textrm{Re}[(\sqrt{2}x_{L}s_{L}^{\nu}c_{L}^{\ell}c^{G})(\lambda_{L}^{\ast}c_{L}^{\nu}c_{R}^{\ell}c^{G})]\cdot F_{LR}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})\,,$
$\approx$ $\quad\quad\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{--}}^{2}}\textrm{Re}[(2x_{L}s_{L}^{\ell})(\lambda_{L}^{\ast})]\cdot[-F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]$
$+\frac{m_{\mu}m_{F^{0}}}{4\pi^{2}m_{\delta^{-}}^{2}}\textrm{Re}[(\sqrt{2}x_{L}s_{L}^{\nu})(\lambda_{L}^{\ast})]\cdot F_{LR}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})\,.$ (65)
Utilizing $s^{\ell}_{L}\approx\sqrt{2}s^{\nu}_{L}$ from Eq. (26), as well as the mass relations $m_{\delta}\equiv m_{\delta^{0}}\approx m_{\delta^{-}}\approx m_{\delta^{--}}$ from Eq. (35) and $m_{F}\equiv m_{F^{\pm}}\approx m_{F^{0}}$ from Eq. (43), we can have the following simple expressions
$a^{\textrm{NP}}_{\mu}\approx\frac{\sqrt{2}m_{\mu}m_{F}}{4\pi^{2}m_{\delta}^{2}}\textrm{Re}[x_{L}\lambda_{L}^{\ast}]s_{L}^{\nu}\cdot[-2F_{LR}^{f,2}(\frac{m_{F}^{2}}{m_{\delta}^{2}})+5F_{LR}^{S,2}(\frac{m_{F}^{2}}{m_{\delta}^{2}})]\,.$ (66)
Based on our calculation we can also reproduce the results of models including only one scalar triplet or one fermion triplet corresponding to Type-II and Type-III seesaw models, respectively. More details can be found in Appendix A.
III.2 Numerical results
[…]
In the left panel of Fig. 2 we impose the $\{\lambda_{L},x_{L}\}$ scheme described in Eq. (II.3) as input parameters. We take $\lambda_{L}=1,\,x_{L}=-0.06$ as benchmark point and it can be seen that $a^{\textrm{NP}}_{\mu}$ predicted by our model can easily cover the central region of $\Delta a_{\mu}$ with $m_{F}\approx 1500\,\textrm{GeV}$. In more details, $a_{\mu}^{F,\,\delta^{--}}$ ($a_{\mu}^{F,\,\delta^{-}}$) indicated by red (blue) line contribute about $85\%$ ($10\%$) to the total $a^{\textrm{NP}}_{\mu}$ in the mass range of $m_{F}$ shown in Fig. 2. Note that according to Eq. (49), only region of $m_{F}\gtrsim 1470\,\textrm{GeV}$ is allowed for $x_{L}=-0.06$ to satisfy $s^{\nu}_{L}\leq 0.017$, i.e. the constraints on SM muon neutrino mixing with heavy neutral lepton as discussed in [39]. We can also see that the decoupling behavior of $a^{\textrm{NP}}_{\mu}$ with increasing $m_{F}$ is clearly manifested. This can be expected from Eq. (II.3) since fixed Yukawa couplings with larger $m_{F}$ would yield smaller mixings $s^{\nu}_{L},s^{\ell}_{L},s^{\ell}_{R}$ and smaller loop function values. More specifically, in the heavy region of $m_{F}$ we can have the following trending behavior of Eq. (66) after utilizing the approximated form of loop functions in Eq. (III.1)
$a^{\textrm{NP}}_{\mu}$ $\approx\frac{\sqrt{2}m_{\mu}m_{F}}{4\pi^{2}m_{\delta}^{2}}(\lambda_{L}x_{L})\,\sqrt{-\sqrt{2}\,x_{L}\frac{v_{\delta}}{m_{F}}}\cdot[-2\,(\frac{1}{4})\,(\frac{m_{F}^{2}}{m_{\delta}^{2}})^{-1}+5\,(-\frac{1}{4})\,(\frac{m_{F}^{2}}{m_{\delta}^{2}})^{-1}]$
$\propto\frac{m_{\mu}}{m_{F}}\lambda_{L}(-x_{L})\sqrt{(-x_{L})\frac{v_{\delta}}{m_{F}}}\,,$ (67)
in which we have taken $x_{L},\lambda_{L}$ to be real numbers to simplify the expression. One can easily see that $a^{\textrm{NP}}_{\mu}$ decreases with fixed $x_{L},\lambda_{L}$ and increasing $m_{F}$. In the right panel of Fig. 2 we impose the $\{\lambda_{L},s^{\nu}_{L}\}$ scheme described in Eq. (II.3) as input parameters and take $\lambda_{L}=1,\,s^{\nu}_{L}=0.017$ as benchmark point. Aside from the observation that $a^{\textrm{NP}}_{\mu}$ can cover the central region of $\Delta a_{\mu}$ with $m_{F}\approx 1500\,\textrm{GeV}$ in this parameter setup, noticeable chiral enhancement is manifested by enhanced $a^{\textrm{NP}}_{\mu}$ when $m_{F}$ increases.
[…]
In Fig. 4 we show $a^{\textrm{NP}}_{\mu}$ on the plane of $m_{F}$ versus $s^{\nu}_{L}$ in which the left (right) panel has $\lambda_{L}=1$ ($\lambda_{L}=0.5$). This figure can be view as the extended illustration of the right panel of Fig. 2 by further allowing $s^{\nu}_{L}$ to vary. The dashed line is $s^{\nu}_{L}=0.017$ with respect to which the region below is allowed. We can see that in the left panel, the region with $s^{\nu}_{L}=0.017$ and $1000\,\textrm{GeV}\lesssim m_{F}\lesssim 2000\,\textrm{GeV}$ can generate $a^{\textrm{NP}}_{\mu}$ within the $1\,\sigma$ range of $\Delta a_{\mu}$. Region with smaller value of $s^{\nu}_{L}$ would require heavier $m_{F}$ and thus more significant chiral enhancement to achieve the same level of $a^{\textrm{NP}}_{\mu}$. The right panel, however, can only generate $a^{\textrm{NP}}_{\mu}$ on the edge of $1\,\sigma$ range of $\Delta a_{\mu}$ with quite heavy fermion mass $m_{F}\approx 5\,\textrm{TeV}$ with $s^{\nu}_{L}=0.017$. We also checked that all Yukawa couplings in our model satisfy the requirement of perturbativity on the shown range of Fig. 4 which can be easily seen through Eq. (II.3).
III.3 Phenomenological discussions
In this section we briefly discuss the phenomenological aspects of our model. Note that the scalar triplet and triplet have been scrutinized in neutrino mass generation mechanisms as Type-II [28, 29, 30, 31, 32, 33] and Type-III [34, 35] seesaw models, respectively. However, in this work we only concentrate on the physics of $(g-2)_{\mu}$ and do not require our model to produce the experimentally suggested texture of neutrino mass matrix. Instead, we require our model parameters to ensure the theoretically predicted neutrino mass to be negligibly small and close to zero. Aside from the neutrino physics, recently it has been found that the Type-II seesaw model can also address the problem of baryon asymmetry of the universe (see e.g. [62, 63, 64, 65]).
For the new physics leptons $F^{0},F^{\pm}$, Eq. (43) indicates that the triplet masses are almost degenerate. Originating from the electroweak gauge interactions, $F^{0},F^{\pm}$ can be singly or pairly produced. The main decay channels of the triplet are $F^{0}\rightarrow\nu h,\nu Z,\ell^{\pm}W^{\mp}$ and $F^{\pm}\rightarrow\ell^{\pm}h,\ell^{\pm}Z,\nu W^{\pm}$. If the mass spectrum satisfies $m_{F}>m_{\delta}$ which is assumed in the analysis of this work, there are additional decay channels $F^{0}\rightarrow\ell^{\pm}\delta^{\mp}$, $F^{-}\rightarrow\ell^{-}\delta^{0},\ell^{-}a^{0},\ell^{+}\delta^{--}$. Similarly, for the new physics scalars $\delta^{0},a^{0},\delta^{-},\delta^{--}$, Eq. (35) indicates that the scalar triplet mass are also nearly degenerate. The scalar triplet can be pairly or singly produced through the $s$-channel mediation of SM gauge bosons $W^{\pm},\,Z$. The fermionic decay channels of the scalar triplet to SM leptons include $\delta^{0}\rightarrow\ell^{+}\ell^{-}$,$a^{0}\rightarrow\ell^{+}\ell^{-}$, $\delta^{-}\rightarrow\nu\ell^{-}$, and $\delta^{--}\rightarrow\ell^{-}\ell^{-}$. The bosonic decay channels of the scalar triplet to SM bosons include $\delta^{0}\rightarrow W^{+}W^{-}$, $\delta^{-}\rightarrow W^{-}Z$, and $\delta^{--}\rightarrow W^{-}W^{-}$, of which the partial decay widths are proportional to $v_{\delta}^{2}$. If the mass spectrum satisfies $m_{\delta}>m_{F}$, there are additional decay channels $\delta^{0}\rightarrow\ell^{\pm}F^{\mp}$, $a^{0}\rightarrow\ell^{\pm}F^{\mp}$, $\delta^{-}\rightarrow F^{0}\ell^{-}$, and $\delta^{--}\rightarrow F^{-}\ell^{-}$. Detailed collider analysis is beyond the scope of this paper, and recent discussions on the prospects of searches for scalar triplet around TeV at future high luminosity LHC and 100 TeV $pp$ collider can be found in [43, 66, 67], respectively. The lepton triplet around TeV can also be accessible at future colliders [68, 69, 70].
IV Conclusions
In this work, we investigated the muon anomalous dipole moment $a_{\mu}$ in a model that extends the SM with a scalar triplet and a lepton triplet. We identify an important ingredient overlooked in previous studies, i.e. the Yukawa interaction involving the SM Higgs doublet and the newly introduced lepton triplet. This interaction is a four-dimensional operator and thus can naturally exist. We show that this Yukawa interaction can not only induce mass mixing between leptons in the SM and the new physics sector, but also provide an additional source of chiral flip to $a_{\mu}$. We find that there is still viable parameter space in this model to explain $\Delta a_{\mu}$, which is different from the observation reported in the existing literature. More specifically, while being consistent with the current data of neutrino mass, electroweak precision measurements and the perturbativity of couplings, our model can provide new physics contribution $a_{\mu}^{\textrm{NP}}$ to cover the central region of $\Delta a_{\mu}$ with new scalar and fermion mass as low as around TeV. This mass scale is allowed by the current collider searches for doubly charged scalars and the lepton triplet, and they can be tested at future high energy and/or high luminosity colliders.
Acknowledgements.
This work was supported by National Key R&D Program of China under grant Nos. 2023YFA1606100. CH is supported by the Sun Yat-Sen University Science Foundation. SH is supported by the research start-up fund of Taiyuan University of Technology. SH would also like to acknowledge the hospitality of Center for High Energy Physics, Peking University, where he spent three months as a visitor. PW acknowledges support from Natural Science Foundation of Jiangsu Province (Grant No. BK20210201), Fundamental Research Funds for the Central Universities, Excellent Scholar Project of Southeast University (Class A), and the Big Data Computing Center of Southeast University.
Appendix A Reproduced results in models including only one scalar triplet or one lepton triplet
If keeping only the scalar triplet or lepton triplet in Section III.1 by turning off relevant interactions, we can reproduce the analytical results of $a_{\mu}$ predicted in Type-II and Type-III seesaw models.
  • By taking $\theta_{L}^{\ell}=\theta_{R}^{\ell}=\theta_{L}^{\nu}=\lambda_{L}=0$, the $a_{\mu}^{F,\,\textrm{total}}$ vanishes automatically and we would reproduce the results in Type-II seesaw model as follows, which are negative and consistent with [24, 61, 71, 27],
    $a_{\mu}^{\ell,\,\textrm{total}}$ $=\frac{m_{\mu}^{2}}{8\pi^{2}}\big{\{}-\frac{2m_{\mu}^{2}(s^{h})^{2}}{m_{h}^{2}v_{h}^{2}}\cdot[F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{h}^{2}})+F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{h}^{2}})]+\frac{2m_{\mu}^{2}(s^{h})^{2}}{m_{\delta^{0}}^{2}v_{h}^{2}}\cdot[F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{0}}^{2}})+F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{0}}^{2}})]$
    $\qquad\quad\,\,\,\,+\frac{4|x_{L}|^{2}}{m_{\delta^{--}}^{2}}\cdot[-F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})+2F_{LL}^{S,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})]\big{\}}\,,$
    $\approx\frac{m_{\mu}^{2}}{8\pi^{2}}\big{\{}-\frac{2m_{\mu}^{2}(s^{h})^{2}}{m_{h}^{2}v_{h}^{2}}[\log(\frac{m_{h}}{m_{\mu}})-\frac{7}{12}]+\frac{2m_{\mu}^{2}(s^{h})^{2}}{m_{\delta^{0}}^{2}v_{h}^{2}}[\log(\frac{m_{\delta^{0}}}{m_{\mu}})-\frac{7}{12}]-\frac{4|x_{L}|^{2}}{3m_{\delta^{--}}^{2}}\big{\}}\,,$
    $a_{\mu}^{\nu,\,\textrm{total}}$ $=\frac{m_{\mu}^{2}|x_{L}|^{2}}{4\pi^{2}m_{\delta^{-}}^{2}}\cdot F_{LL}^{S,3}(\frac{m_{\mu}^{2}}{m_{\delta^{-}}^{2}})\approx-\frac{m_{\mu}^{2}|x_{L}|^{2}}{48\pi^{2}m_{\delta^{-}}^{2}}\,.$ (68)
  • By taking $\theta^{h}=\theta^{a}=\theta^{G}=x_{L}=\lambda_{L}=0$, the $a_{\mu}^{\nu,\,\textrm{total}}$ vanishes automatically and we would reproduce the results in Type-III seesaw model as follows
    $a_{\mu}^{\ell,\,\textrm{total}}$ $=\frac{m_{\mu}^{4}}{4\pi^{2}m_{h}^{2}v_{h}^{2}}[(c_{L}^{\ell})^{4}-1]\cdot[F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{h}^{2}})+F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{h}^{2}})]$
    $\approx-\frac{m_{\mu}^{4}(s_{L}^{\ell})^{2}}{2\pi^{2}m_{h}^{2}v_{h}^{2}}(\log(\frac{m_{h}}{m_{\mu}})-\frac{7}{12})\,,$
    $a_{\mu}^{F,\,\textrm{total}}$ $=\frac{m_{\mu}^{2}m_{F^{\pm}}^{2}}{8\pi^{2}m_{h}^{2}v_{h}^{2}}(s_{L}^{\ell})^{2}(c_{L}^{\ell})^{2}[(1+\frac{m_{\mu}^{2}}{m_{F^{\pm}}^{2}})F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{h}^{2}})+2F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{h}^{2}})]$
    $\approx\frac{m_{\mu}^{2}m_{F^{\pm}}^{2}}{8\pi^{2}m_{h}^{2}v_{h}^{2}}(s_{L}^{\ell})^{2}(c_{L}^{\ell})^{2}[F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{h}^{2}})+2F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{h}^{2}})]\,.$ (69)